# Bayesian Optimization for Non-Convex Two-Stage Stochastic Optimization Problems

### v1

@@ Abstract
Bayesian optimization is a sample-efficient method for solving expensive, black-box optimization problems. Stochastic programming concerns optimization under uncertainty where, typically, average performance is the quantity of interest. In the first stage of a two-stage problem, here-and-now decisions must be made in the face of this uncertainty, while in the second stage, wait-and-see decisions are made after the uncertainty has been resolved. Many methods in stochastic programming assume that the objective is cheap to evaluate and linear or convex. In this work, we apply Bayesian optimization to solve non-convex, two-stage stochastic programs which are expensive to evaluate. We formulate a knowledge-gradient-based acquisition function to jointly optimize the first- and second-stage variables, establish a guarantee of asymptotic consistency and provide a computationally efficient approximation. We demonstrate comparable empirical results to an alternative we formulate which alternates its focus between the two variable types, and superior empirical results over the standard, naïve, two-step benchmark. We show that differences in the dimension and length scales between the variable types can lead to inefficiencies of the two-step algorithm, while the joint and alternating acquisition functions perform well in all problems tested. Experiments are conducted on both synthetic and real-world examples.
@@ 1 Introduction
Many optimization problems involve decisions that must be made without full information. This uncertainty can be handled in several ways depending on the risk appetite of the decision maker, ranging from optimizing for the worst-case outcome, to optimizing for average performance. The Bayesian optimization community refers to all these approaches as robust optimization problems, however, the wider optimization community usually reserve the term robust for worst-case optimization. Optimization of the average case is the default for stochastic programming (Birge & Louveaux, 2011; Shapiro et al., 2021).
The notion of a two-stage decision process was first introduced to stochastic programming by Dantzig (1955), who considered linear programs. In contrast, we consider non-convex problems, but the setting is otherwise the same.
We wish to maximize some function $h({\bm{x}},{\mathbf{y}},{\mathbf{u}})$, whose value is dependent on a random variable ${\mathbf{u}}$ representing the uncertainty in the system. We wish to optimize this in the sense of expected performance, $\mathbb{E}_{\mathbf{u}}[h({\bm{x}},{\mathbf{y}},{\mathbf{u}})]$. In the first stage of the optimization, we choose the here-and-now variables or fixed design, ${\bm{x}}$, before the value taken by the uncertain variable ${\mathbf{u}}$ is known. In the second stage, we choose the wait-and-see or adjustable variables, ${\mathbf{y}}$, which can depend on the value of ${\mathbf{u}}$. Thus, the aim is to learn an optimal vector ${\bm{x}}$ and function ${\mathbf{y}}={\bm{g}}({\mathbf{u}})$ to maximize the expected objective. This is illustrated in Figure 1.
@@
Figure 1: A contour plot for a non-convex two-stage stochastic optimization problem at three different values of the fixed design, ${\bm{x}}$. The optimal second-stage decision, ${\bm{y}}={\bm{g}}({\bm{u}})$, is shown by the solid blue line. The aim is to find the fixed design with the best objective value after taking the expectation over the environmental variable.
@@
Non-convex, two-stage stochastic programs arise throughout operations research and engineering, including in the design of wind farms (Chen et al., 2022) and the management of power stations (Phan & Ghosh, 2014). Similar problems are popular in optimal control, specifically in control co-design, where the objective is to jointly optimize a fixed plant design and a control signal in the presence of a feedback loop. Example applications include hybrid electric vehicles (Yan et al., 2020; Xun et al., 2022), batteries (Cui & Wang, 2021) and wind turbines (Cui et al., 2021).
@@
In many of these problems, evaluating the objective is expensive. For example, it may require running time-consuming fluid dynamics or finite element analysis simulations, or be a physical experiment that not only takes time but consumes costly resources. It therefore pays to put more computational effort into reducing the number of evaluations required compared with a typical stochastic approximation approach with numerical gradients (Shapiro et al., 2021, Section 8.2). Bayesian optimization is a sample-efficient, global optimization technique designed to tackle such problems. A probabilistic model of the objective – usually a Gaussian process – is maintained, and an acquisition function is used to measure the value of collecting a sample at a proposed input.
@@
We can naïvely apply Bayesian optimization to solve two-stage stochastic programs, by first optimizing the wait-and-see control policy ${\mathbf{y}}={\bm{g}}({\mathbf{u}})$ for an arbitrary value of the fixed design ${\bm{x}}$, then switching to optimize the fixed design for the control policy found in the first phase. The two phases could be alternated until convergence. However, each phase requires a new initial sample, waiting for each step to converge is wasteful and slow, and the iteration is not guaranteed to converge to the global optimum. There is therefore potential to improve greatly upon this by jointly optimizing both sets of variables.
@@
In this paper, we formulate the knowledge gradient acquisition function for two-stage stochastic optimization problems without coupled constraints. We establish asymptotic consistency of the estimators of the maximum of the expected performance, $\mathbb{E}_{\mathbf{u}}[h({\bm{x}}^{*n},{\bm{g}}^{*n}({\mathbf{u}}),{\mathbf{u}})]$, associated with the fixed design ${\bm{x}}^{*n}$ and control policy ${\bm{g}}^{*n}$ which would be recommended with this acquisition function after $n$ iterations, under the assumption that the function $h$ is drawn from a Gaussian process with known mean and covariance function. Further, we provide a computationally efficient approximation scheme to optimize this acquisition function and present an alternative algorithm which alternates its focus between the here-and-now and wait-and-see variables, making fewer approximations. We demonstrate empirically that both of these perform similarly on a suite of synthetic and real-world problems, with and without observation noise, and significantly outperform the standard ‘two-step’ algorithm which alternates between fully optimizing ${\bm{x}}$ and fully optimizing ${\bm{y}}={\bm{g}}({\bm{u}})$. Increasing the dimension of ${\mathbb{X}}$ and ${\mathbb{Y}}$ and decreasing the length scale of $h$ makes different aspects of the problem harder and we investigate the effect of this on the three algorithms.
@@ 1.1 Problem statement
Let ${\mathbb{X}}\subset\mathbb{R}^{d_{x}}$, ${\mathbb{Y}}\subset\mathbb{R}^{d_{y}}$ and ${\mathbb{U}}\subset\mathbb{R}^{d_{u}}$ denote the spaces of fixed, adjustable and environmental variables respectively. Let $h:{\mathbb{X}}\times{\mathbb{Y}}\times{\mathbb{U}}\to\mathbb{R}$ be an unknown function and suppose that we can make expensive observations of $h$ at points of our choosing $({\bm{x}},{\bm{y}},{\bm{u}})\in{\mathbb{X}}\times{\mathbb{Y}}\times{\mathbb{U}}$, possibly corrupted by Gaussian noise,
@@
This paper tackles the problem of finding fixed parameters ${\bm{x}}\in{\mathbb{X}}$ and a control policy ${\bm{g}}:{\mathbb{U}}\to{\mathbb{Y}}$ which maximize the expectation of $h$ over the uncertain environment, ${\mathbf{u}}$. That is, it solves
@@
Note that the problem can also be phrased in the classical manner as a stochastic bi-level optimization, by moving the optimization of ${\bm{g}}$ inside the expectation,
@@
Any constraints on ${\bm{x}}$ and ${\bm{y}}$ are assumed part of the definition of ${\mathbb{X}}$ and ${\mathbb{Y}}$. We do not consider coupled constraints where the space ${\mathbb{Y}}$ of admissible ${\bm{y}}$ depends on ${\bm{x}}$, nor when either of ${\mathbb{X}}$ and ${\mathbb{Y}}$ depend on ${\mathbf{u}}$.
Throughout this paper, we use an italic font, ${\bm{u}}$, to represent deterministic quantities and a straight font, ${\mathbf{u}}$, for random vectors. The exception to this are Greek letters, such as the observation noise $\varepsilon$, and the Gaussian process, $f$, to be introduced later.
@@ 1.2 Related work
Non-convex, two-stage stochastic optimization problems are a class of uncertain bi-level optimization problem where the inner and outer objectives are identical. Many more sources of uncertainty are possible in this setup beyond the uncertainty from ${\mathbf{u}}$, arising when the first and second level decision makers do not know how each other will react (Beck et al., 2023). Bayesian optimization with the UCB acquisition function has been applied to bi-level problems in (Fu et al., 2024), while a surrogate-assisted evolutionary algorithm was employed in (Sinha & Shaikh, 2022) for bi-level problems.
@@
Salomon (2019) explore non-convex, two-stage optimization problems under uncertainty under the name active robust optimization. Various risk measures are considered, including expectation, worst-case performance, value-at-risk, and target-based. Both single-objective and multi-objective problems are considered, and evolutionary algorithms are proposed for the multi-objective case.
@@
Finally, we come to Bayesian optimization. To the best of the authors’ knowledge, two-stage stochastic optimization problems have not yet been tackled with Bayesian optimization. However, the problem is closely related to both standard stochastic optimization (the first-stage decision) and multi-task optimization (the second-stage decision).
@@
Robust Bayesian optimization aims to solve $\max_{{\bm{x}}\in{\mathbb{X}}}\mathbb{E}_{\mathbf{u}}[h({\bm{x}},{\mathbf{u}})]$ and has received a lot of attention, with authors proposing acquisition strategies based on expected improvement Williams et al. (2000); Groot et al. (2010) and knowledge gradient (Pearce & Branke, 2017; Pearce et al., 2022; Toscano-Palmerin & Frazier, 2022) to name a few. The related problem of robustness to perturbed or noisy inputs has also received much attention (Nogueira et al., 2016; Beland & Nair, 2017; Oliveira et al., 2019; Fröhlich et al., 2020; Le & Branke, 2021; 2024), as has the study of other risk measures such as value-at-risk (Cakmak et al., 2020; Daulton et al., 2022) and mean-variance trade-offs (Qing et al., 2022). Generally, it is assumed that the distribution of the environmental variables is known, but recent work exists to infer this distribution (Huang et al., 2024).
@@
In multi-task Bayesian optimization, the goal is to solve a finite set, or continuum of optimization problems, parameterized by ${\bm{u}}$. This amounts to finding the control policy that optimizes the inner part of Equation 3, $\max_{{\bm{g}}:{\mathbb{U}}\to{\mathbb{Y}}}\mathbb{E}_{\mathbf{u}}[h({\bm{g}}({\mathbf{u}}),{\mathbf{u}})]=\mathbb{E}_{\mathbf{u}}[\max_{{\mathbf{y}}\in{\mathbb{Y}}}h({\mathbf{y}},{\mathbf{u}})]$. Acquisition functions based on entropy (Swersky et al., 2013), expected improvement (Tesch et al., 2011; Ginsbourger et al., 2014) and knowledge gradient (Pearce & Branke, 2018; Ding et al., 2022) have been proposed previously.
@@ 2.1 Gaussian processes
When used to model an unknown function, the prior mean and covariance functions of a GP should be chosen to fit the function being modelled. Popular choices for the covariance function include the squared exponential kernel and Matérn family of kernels. These are parameterized by a length scale for each dimension and an output scale. The mean function is often chosen to be zero or a constant. The length scales, output scale and constant mean, together with the variance of the observation noise are known as hyperparameters of the GP. They can be fitted by placing prior distributions on the hyperparameters and using maximum a posteriori (MAP) estimates. Rasmussen & Williams (2006) provide an excellent introduction to Gaussian processes for machine learning.
@@ 2.2 Bayesian optimization
The role of $\alpha$ is to balance exploration of regions of the input space with high uncertainty, and exploitation of regions already known to contain good values. Most commonly, the Bayesian model consists of a Gaussian process (GP) prior, along with an observation model with additive Gaussian observation noise such as Equation 1.
@@ 2.3 Knowledge gradient
A popular acquisition function capable of handling observation noise and partial information is the knowledge gradient (Frazier et al., 2009; Scott et al., 2011). The grounding principle is that at the end of the data collection process, the recommended solution will be the maximum of the posterior mean. Write $\mathbb{E}_{n}$ for expectation conditional on the $n$ observations made during the optimization, and let $\mu^{n}({\bm{x}})=\mathbb{E}_{n}[f({\bm{x}})]$ be the posterior mean of the GP conditional on these observations. The final recommendation is
${\bm{x}}^{*n}=\operatorname*{arg\,max}_{{\bm{x}}\in{\mathbb{X}}}\mu^{n}({\bm{x}}).$ (4)
The knowledge gradient acquisition function is then constructed to be one-step Bayes-optimal. That is, maximizing it corresponds to maximizing $\mathbb{E}_{n}[f({\mathbf{x}}^{*(n+1)})]$, where here the expectation is taken over both $f$ and the next observation. Let ${\textnormal{v}}^{n+1}=f({\bm{x}}^{n+1})+\varepsilon^{n+1}$ be the random variable representing the next observation. The knowledge gradient is given by
$\alpha_{\text{KG}}^{n}({\bm{x}})=\mathbb{E}_{{\textnormal{v}}^{n+1}}\biggl{[}\max_{{\bm{x}}^{\prime}\in{\mathbb{X}}}\mu^{n+1}({\bm{x}}^{\prime})\,\bigg{|}\,{\bm{x}}^{n+1}={\bm{x}}\biggr{]}-\max_{{\bm{x}}^{\prime}\in{\mathbb{X}}}\mu^{n}({\bm{x}}^{\prime}).$ (5)

### v2

@@ Abstract
Bayesian optimization is a sample-efficient method for solving expensive, black-box optimization problems. Stochastic programming concerns optimization under uncertainty where, typically, average performance is the quantity of interest. In the first stage of a two-stage problem, here-and-now decisions must be made in the face of uncertainty, while in the second stage, wait-and-see decisions are made after the uncertainty has been resolved. Many methods in stochastic programming assume that the objective is cheap to evaluate and linear or convex. We apply Bayesian optimization to solve non-convex, two-stage stochastic programs which are black-box and expensive to evaluate as, for example, is often the case with simulation objectives. We formulate a knowledge-gradient-based acquisition function to jointly optimize the first- and second-stage variables, establish a guarantee of asymptotic consistency, and provide a computationally efficient approximation. We demonstrate comparable empirical results to an alternative we formulate with fewer approximations, which alternates its focus between the two variable types, and superior empirical results over the state of the art and the standard, naïve, two-step benchmark.
@@ 1 Introduction
Many optimization problems involve decisions that must be made with incomplete information. This uncertainty can be handled in several ways depending on the risk appetite of the decision maker, ranging from optimizing for the worst-case outcome, to optimizing for average performance. The Bayesian optimization community refers to all these approaches as robust optimization problems; however, the operations research community mostly reserves the term robust optimization for worst-case settings. Optimization of the average case is the default for stochastic programming (Birge & Louveaux, 2011; Shapiro et al., 2021).
The notion of a two-stage decision process was first introduced to stochastic programming by Dantzig (1955), who considered linear programs. In contrast, we consider non-convex problems but the setting is the same. We wish to maximize some objective in expectation over a stochastic environment. In the first stage of the optimization, we choose the here-and-now variables or fixed design before the uncertainty in the environment is resolved. In the second stage, we choose the wait-and-see or adjustable variables once the specific environment has been revealed. The aim is to learn both an optimal fixed design and a control policy mapping environments to adjustable variables to maximize the expected objective. This is illustrated in Figure 1.
@@
Figure 1: A contour plot for a non-convex two-stage stochastic optimization problem at three different values of the fixed design, ${\bm{x}}$. The solid blue line shows the optimal second-stage decision, ${\bm{y}}={\bm{g}}({\bm{u}})$. The aim is to find the fixed design with the best objective value after taking the expectation over the environmental variable.
In many cases, including cases from simulation optimization, the objective is a black box. That is, gradients are typically not available and the simulation is so complicated that there are no insights exploitable to make the optimization easier. The only available tool is to evaluate the objective at specific inputs. Furthermore, these evaluations are often expensive. For example, they may require running time-consuming fluid dynamics or finite element analysis simulations, or be a physical experiment that not only takes time but consumes costly resources. This variation has so far received relatively little attention in the literature and to the best of the authors’ knowledge, (Xie et al., 2021) is the only other work to address it.
@@
Examples of these problems arise throughout operations research and engineering, including in the design of wind farms (Chen et al., 2022) and the management of power stations (Phan & Ghosh, 2014). Similar problems are popular in optimal control, specifically in control co-design, where the objective is to jointly optimize a fixed plant design and a control signal in the presence of a feedback loop. Example applications include hybrid electric vehicles (Yan et al., 2020; Xun et al., 2022), batteries (Cui & Wang, 2021) and wind turbines (Cui et al., 2021).
@@
Since evaluations of the objective are expensive, it pays to put more computational effort into reducing the number of evaluations required compared with a typical stochastic approximation approach with numerical gradients (Shapiro et al., 2021, Section 8.2). Bayesian optimization is a sample-efficient, global optimization technique designed to tackle such problems. A probabilistic model of the objective – usually a Gaussian process – is maintained, and an acquisition function is used to measure the value of collecting a sample at a proposed input.
@@
We can naïvely apply Bayesian optimization to solve two-stage stochastic programs, by first optimizing the wait-and-see control policy for an arbitrary value of the fixed design, then switching to optimize the fixed design for the control policy found in the first phase. The two phases could be alternated until convergence. However, each phase requires a new initial sample, waiting for each step to converge is wasteful and slow, and the iteration is not guaranteed to converge to the global optimum. There is therefore potential to improve greatly upon this by jointly optimizing both sets of variables.
@@
In this work, we formulate the knowledge gradient acquisition function for two-stage stochastic optimization problems without coupled constraints. We establish asymptotic consistency of the estimators of the maximum of the expected performance associated with the fixed design and control policy which would be recommended with this acquisition function at each iteration, under the assumption that the objective is drawn from a Gaussian process with known mean and covariance function. Further, we provide a computationally efficient approximation scheme to optimize this acquisition function and present an alternative algorithm that alternates its focus between the here-and-now and wait-and-see variables, making fewer approximations. We demonstrate empirically that both of these perform similarly on a suite of synthetic and real-world problems, with and without observation noise, and significantly outperform the standard ‘two-step’ algorithm which alternates between fully optimizing the fixed design and fully optimizing the control policy. Increasing the dimension and decreasing the intrinsic length scales of the fixed and adjustable variables makes different aspects of the problem harder and we investigate the effect of this on the three algorithms. Finally, we establish superior performance of the joint algorithm over the state of the art on a real-world supply chain example.
@@ 1.1 Problem statement
Let ${\mathbb{X}}\subset\mathbb{R}^{d_{x}}$, ${\mathbb{Y}}\subset\mathbb{R}^{d_{y}}$ and ${\mathbb{U}}\subset\mathbb{R}^{d_{u}}$ denote the spaces of fixed, adjustable and environmental variables respectively. Let $h:{\mathbb{X}}\times{\mathbb{Y}}\times{\mathbb{U}}\to\mathbb{R}$ be an unknown function, say given by a simulation, and suppose that we can make expensive observations of $h$ at points $({\bm{x}},{\bm{y}},{\bm{u}})\in{\mathbb{X}}\times{\mathbb{Y}}\times{\mathbb{U}}$ of our choosing, possibly corrupted by Gaussian noise,
@@
This work tackles the problem of finding fixed parameters ${\bm{x}}\in{\mathbb{X}}$ and a control policy ${\bm{g}}:{\mathbb{U}}\to{\mathbb{Y}}$ for the adjustable parameters which maximize the expectation of $h$ over the uncertain environment, ${\mathbf{u}}$. That is, it solves
@@
Note that the problem can also be phrased classically as a stochastic bi-level optimization problem, by moving the optimization of ${\bm{g}}$ inside the expectation,
@@
Any constraints on ${\bm{x}}$ and ${\bm{y}}$ are assumed part of the definition of ${\mathbb{X}}$ and ${\mathbb{Y}}$. Except for the supply chain problem in Section 4.4, we do not consider coupled constraints where the space ${\mathbb{Y}}$ of admissible ${\bm{y}}$ depends on ${\bm{x}}$, nor when either of ${\mathbb{X}}$ and ${\mathbb{Y}}$ depend on ${\mathbf{u}}$.
Throughout this work, we use an italic font, ${\bm{u}}$, to represent deterministic quantities and a straight font, ${\mathbf{u}}$, for random vectors. The exception to this are Greek letters, such as the observation noise $\varepsilon$, and the Gaussian process, $f$, to be introduced later.
@@ 1.2 Related work
Non-convex, two-stage stochastic optimization problems are a class of uncertain bi-level optimization where the inner and outer objectives are identical. Many more sources of uncertainty are possible in this setup beyond the uncertainty from ${\mathbf{u}}$, arising when the first and second level decision makers do not know how each other will react (Beck et al., 2023). Bayesian optimization with the Upper Confidence Bound (UCB) acquisition function has been applied to bi-level problems in (Fu et al., 2024), while a surrogate-assisted evolutionary algorithm was employed in (Sinha & Shaikh, 2022) for bi-level problems.
@@
Salomon (2019) explore non-convex, two-stage optimization problems under uncertainty as active robust optimization. Various risk measures are considered, including expectation, worst-case performance, value-at-risk, and target-based. Both single-objective and multi-objective problems are considered, and evolutionary algorithms are proposed for the multi-objective case.
@@
Finally, two-stage stochastic optimization problems are closely related to standard stochastic optimization (the first-stage decision) and multi-task optimization (the second-stage decision), and Bayesian optimization has been widely applied in both cases.
@@
Stochastic, or robust, Bayesian optimization aims to solve $\max_{{\bm{x}}\in{\mathbb{X}}}\mathbb{E}_{\mathbf{u}}[h({\bm{x}},{\mathbf{u}})]$ and has received a lot of attention, with authors proposing acquisition strategies based on expected improvement (Williams et al., 2000; Groot et al., 2010) and knowledge gradient (Pearce & Branke, 2017; Pearce et al., 2022; Toscano-Palmerin & Frazier, 2022) to name a few. The related problem of robustness to perturbed or noisy inputs has also received much attention (Nogueira et al., 2016; Beland & Nair, 2017; Oliveira et al., 2019; Fröhlich et al., 2020; Le & Branke, 2021; 2024), as has the study of other risk measures such as value-at-risk (Cakmak et al., 2020; Daulton et al., 2022) and mean-variance trade-offs (Qing et al., 2022). Generally, it is assumed that the distribution of the environmental variables is known, but recent work exists to infer this distribution (Huang et al., 2024).
@@
In multi-task Bayesian optimization, the goal is to solve a finite set or continuum of optimization problems, $\max_{\bm{y}}h({\bm{y}},{\bm{u}})$, parameterized by ${\bm{u}}$ as in Equation 3b. Acquisition functions based on entropy (Swersky et al., 2013), expected improvement (Tesch et al., 2011; Ginsbourger et al., 2014) and knowledge gradient (Pearce & Branke, 2018; Ding et al., 2022) have been proposed previously.
The closest work to our own is that of Xie et al. (2021) who tackle expensive, non-convex, two-stage stochastic optimization problems by adopting a nested approach. They use a Gaussian process surrogate model both on the inner level to optimize the control policy ${\bm{y}}={\bm{g}}({\bm{x}})$ and on the outer level to optimize the fixed design ${\bm{x}}$. The problem setting differs slightly from the one considered in Equation 2 in that Xie et al. (2021) only aim to find the optimal first-stage decision ${\bm{x}}$, while we also seek to return an optimal control policy ${\bm{y}}={\bm{g}}({\bm{x}})$. Nonetheless, the setting is sufficiently similar for us to provide a numerical comparison in Section 4.
@@ 2.1 Gaussian processes
When used to model an unknown function, the prior mean and covariance functions of a GP should be chosen to fit the function being modeled. Popular choices for the covariance function include the squared exponential kernel and Matérn family of kernels. These are parameterized by a length scale for each dimension and an output scale. The mean function is often chosen to be zero or a constant. The length scales, output scale, and constant mean, together with the variance of the observation noise are known as hyperparameters of the GP. They can be fitted by placing prior distributions on the hyperparameters and using maximum a posteriori (MAP) estimates. Rasmussen & Williams (2006) provide an excellent introduction to Gaussian processes for machine learning.
@@ 2.2 Bayesian optimization
The role of $\alpha$ is to balance exploration of regions of the input space with high uncertainty, and exploitation of regions already known to contain good values. Most commonly, the Bayesian model consists of a GP prior, along with an observation model with additive Gaussian noise such as Equation 1.
@@ 2.3 Knowledge gradient
A popular acquisition function capable of handling observation noise and partial information is the knowledge gradient (KG) (Frazier et al., 2009; Scott et al., 2011). The grounding principle is that at the end of the data collection process, the recommended solution will be the maximum of the posterior mean. Write $\mu^{n}({\bm{x}})=\mathbb{E}[f({\bm{x}})\,|\,\mathcal{D}^{n}]$ for the posterior mean of the GP conditional on the $n$ observations made during the optimization. The final recommendation is
${\bm{x}}^{*n}\in\operatorname*{arg\,max}_{{\bm{x}}\in{\mathbb{X}}}\mu^{n}({\bm{x}}).$ (4)
The knowledge gradient acquisition function is then constructed to be one-step Bayes-optimal. That is, maximizing it corresponds to maximizing the expected value at the recommended solution, $\mathbb{E}[f({\mathbf{x}}^{*(n+1)})\,|\,\mathcal{D}^{n}]$, where here the expectation is taken over both $f$ and the next observation. Let ${\textnormal{v}}^{n+1}=f({\bm{x}}^{n+1})+\varepsilon^{n+1}$ be the random variable representing the next observation. The knowledge gradient is given by
$\alpha_{\text{KG}}^{n}({\bm{x}})=\mathbb{E}_{{\textnormal{v}}^{n+1}}\biggl{[}\max_{{\mathbf{x}}^{\prime}\in{\mathbb{X}}}\mu^{n+1}({\mathbf{x}}^{\prime})\,\bigg{|}\,{\bm{x}}^{n+1}={\bm{x}}\biggr{]}-\max_{{\bm{x}}^{\prime}\in{\mathbb{X}}}\mu^{n}({\bm{x}}^{\prime}).$ (5)